# Deep Legendre Transform

### v1

@@ Abstract
We introduce a novel deep learning algorithm for computing convex conjugates of differentiable convex functions, a fundamental operation in convex analysis with various applications in different fields such as optimization, control theory, physics and economics. While traditional numerical methods suffer from the curse of dimensionality and become computationally intractable in high dimensions, more recent neural network-based approaches scale better, but have mostly been studied with the aim of solving optimal transport problems and require the solution of complicated optimization or max-min problems. Using an implicit Fenchel formulation of convex conjugation, our approach facilitates an efficient gradient-based framework for the minimization of approximation errors and, as a byproduct, also provides a posteriori error estimates for the approximation quality. Numerical experiments demonstrate our method’s ability to deliver accurate results across different high-dimensional examples. Moreover, by employing symbolic regression with Kolmogorov–Arnold networks, it is able to obtain the exact convex conjugates of specific convex functions. Code is available at https://github.com/lexmar07/Deep-Legendre-Transform
@@ 1 Introduction
In this paper, we introduce a deep learning method for computing convex conjugates, which play an important role in different fields such as convex analysis, optimization, control theory, physics, and economics.
The modern formulation of convex conjugation, also called Legendre–Fenchel transformation, was introduced by Fenchel49  [^1] The Legendre transform—a fundamental operation for switching between dual formulations and variables—has broad utility across thermodynamics, mechanics, optimization, and economics. Key applications include: switching between thermodynamic potentials by exchanging extensive variables (entropy, volume) for intensive conjugates (temperature, pressure) in physics; constructing Moreau envelopes in variational analysis; deriving indirect utility and profit functions in economics; and computing convex potentials in optimal transport. Recent work has further demonstrated applications in approximate dynamic programming kolarijani2023fast and self-concordant smoothing adeoye2023self. This paper introduces the Deep Legendre Transform algorithm, establishes its theoretical foundations, and demonstrates its application to optimal-control problems governed by Hamilton–Jacobi equations. and defines the convex conjugate of a function $f\colon C\to\mathbb{R}$ on a nonempty subset $C\subseteq\mathbb{R}^{d}$ by
@@
for all $y\in D=\nabla f(C)$. The right side of (1.2), called Legendre transform, was used by Leg when studying the minimal surface problem. Note that for it to be well-defined, the gradient map $\nabla f\colon C\to D$ does not have to be invertible. In fact, it can easily be shown that if $f$ is differentiable and convex, the right side of (1.1) is maximized by any $x\in C$ satisfying $\nabla f(x)=y$. So the pre-image $(\nabla f)^{-1}(y)$ in (1.2) can be substituted by any point $x\in(\nabla f)^{-1}(y)$.
@@
based on a finite subset $\mathcal{X}\subseteq C$. Of course, for (1.3) to provide a satisfactory approximation to $f^{*}$, the subsets $\mathcal{X}$ and $\mathcal{Y}$ need to be sufficiently dense. Typically, $\mathcal{X}$ and $\mathcal{Y}$ are chosen as grids $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{d}$ and $\mathcal{Y}=\mathcal{Y}_{1}\times\dots\times\mathcal{Y}_{d}$ of similar size. For instance, for regular grids of the form $\mathcal{X}=\mathcal{Y}=I_{N}^{d}=\{\frac{1}{N},\frac{2}{N},\dots,1\}^{d}$, conditions under which the discrete Legendre–Fenchel transform (1.3) converges to $f^{*}$ have been given by corrias1996fast. In this case, a naive computation of (1.3) for all $y\in\mathcal{Y}$, has complexity $O(N^{2d})$. For the case where NN is a power of 22, brenier and Noullez have derived fast algorithms to compute (1.3), similar in spirit to the fast Fourier transform of CoTu, with complexity $O((N\log N)^{d})$. By first passing to convex hulls, Lucet1997 has been able to design an algorithm with complexity $O(|\mathcal{X}|+|\mathcal{Y}|)$, which for $\mathcal{X}=\mathcal{Y}=I^{d}_{N}$ means $O(N^{d})$. But since this is still exponential in $d$, it essentially limits the method to dimensions $d=1,2,3$ and $4.$
A different approach, going back to a tweet by Pey, uses the fact that if $C\subseteq\mathbb{R}^{d}$ is bounded, $f^{*}$ can be approximated [^2] The entropic log-sum-exp smoothing of a max can be traced back to the Gibbs variational principle; brought into modern convex-optimisation analysis by Nesterov2005. with the softmax convex conjugate
@@
Approximating the convolution $Q_{\varepsilon}(f)*G_{\varepsilon}$ with a sum over a finite set $\mathcal{X}\subseteq C$ yields a GPU-friendly algorithm with complexity $O(|\mathcal{X}|\log|\mathcal{X}|)$; see BlondelRouletGoogle24 for details. But since it is grid-based, it still only gives accurate results in low-dimensional settings.
LucetBauschkeHeinzTrienis2009 have explored the computation of convex conjugates of bivariate piecewise linear quadratic convex functions, and HaqueLucet18 have proposed an algorithm for such functions with linear worst-case complexity. But the scalability to higher dimensions remains a challenge.
Recently, numerical methods for convex conjugation have also been studied in the machine learning literature as a tool to solve the Kantorovich dual of optimal transport problems and develop efficient Wasserstein GANs. While their performance has mostly been tested to see how well they helped to solve transport problems, promising results have been obtained by training neural networks to predict an optimal $x$ in the Legendre–Fenchel transform (1.1). E.g. Hoang2019 and Tag2019 predict $x$ from $y$ with a a neural network $h_{\vartheta}(y)$, where $h_{\vartheta}$ is maximizing the right side of (1.1). Mak2020 do the same using predictors of the form $h_{\vartheta}=\nabla\ell_{\vartheta}$ for an ICNN [^3] Input Convex Neural Networks; see Amos2017 $\ell_{\vartheta}$. Korotin2019; Korotin2021 and Amos2023 have proposed to predict an optimal $x$ in (1.1) with a neural network $h_{\vartheta}$ by minimizing the first order condition
@@
$g_{\theta}(\nabla f(x))\approx\left<x,\nabla f(x)\right>-f(x)\quad\mbox{for all }x\in C,$
which can be formulated as a learning problem of the form
@@
    If $g_{\theta}$ is chosen as an ICNN or ResNet the approach readily scales to high-dimensional setups, and if combined with an ICNN as approximator, it is guaranteed to output a convex approximation to the true Legendre transform, which has to be convex.
@@
    Specifying $g_{\theta}$ as a KAN [^4] Kolmogorov–Arnold Network; see KAN makes it possible to obtain exact solutions for certain convex functions.
@@
(i) is a theoretical advantage compared to existing approaches and, as illustrated in Section 5 below, yields good numerical results, in fact, limited only by the used NN architecture/computing power available. We demonstrate that the method achieves state-of-the-art performance by comparing it to the direct learning of the convex conjugate ((i.e., using f^* as a target function in cases where it is available in closed form).
(ii) ICNNs have been introduced by Amos2017, and it has been shown by chen2019optimal that they can approximate any convex function arbitrarily well on compacts. With neural networks as approximators, the minimization problem (1.5) can numerically be solved in high dimensions with a stochastic gradient descent algorithm. The choice of ICNNs ensures that the approximation is convex. This regularizes the learning problem and helps against overfitting. Moreover, it is helpful in various downstream tasks, such as optimization, control, or optimal transport, where convex conjugation can be utilized as a powerful tool to find optimal solutions. Alternatively, we use ResNet (he2016deep), which in many multidimensional cases performs better than ICNN, although convexity is not guaranteed.
(iii) KANs have recently been introduced by KAN. They are particularly powerful for approximation tasks requiring high precision. But in addition, they can also be used for symbolic regression to recover functions that have a closed form expressions in terms of special functions such as polynomials, $\exp$, $\log$, $\sin$ and $\cos$. In Section 5.5 below, we show that KANs can compute Legendre transforms of certain convex functions exactly.
(iv) As a byproduct, our method yields a posteriori $L^{2}$-approximation guarantees. Indeed, by testing a candidate approximating function $g\colon C\to\mathbb{R}$ against the implicit Legendre transform (1.4) on a test set $\mathcal{X}_{\rm test}\subseteq C$ sampled from a distribution $\mu$, it is possible to derive estimates of the $L^{2}$-approximation error $\left\|g-f^{*}\right\|^{2}_{L^{2}(D,\nu)}$ with respect to the push-forward $\nu=\mu\circ(\nabla f)^{-1}$ of $\mu$ under the gradient map $\nabla f\colon C\to D$. This is valuable beyond the framework of this paper since it makes it possible to evaluate the performance of any numerical method for convex conjugation.
@@
The remainder of the paper is organized as follows: In Section 2, we explain how the Legendre transform can be formulated as a machine learning problem. In Section 3, we derive estimates of the $L^{2}$-approximation error of any candidate approximating function of $f^{*}$ with respect to a given probability measure $\nu$ on the image $D=\nabla f(C)$ of $C$ under the gradient map $\nabla f$. In Section 4, we learn a generalized inverse mapping of $\nabla f$ that improves the performance of the method if $\nabla f$ is heavily distorting $C$ and is needed to estimate the $L^{2}$-approximation error with respect to a desired target distribution $\nu$ on $D$. In Section 5, we report the results of different numerical experiments.
@@ 2 Convex conjugation as a machine learning problem
Let $f\colon C\to\mathbb{R}$ be a differentiable convex function defined on an open convex set $C\subseteq\mathbb{R}^{d}$. Our goal is to approximate the convex conjugate $f^{*}\colon\mathbb{R}^{d}\to(-\infty,\infty]$ on the image $D=\nabla f(C)$ of $C$ under the gradient map $\nabla f$.
By the Fenchel–Young inequality555see e.g., Fenchel49 or Rockafellar70, one has
@@
which makes it possible to formulate the numerical approximation of $f^{*}$ as a machine learning problem. More precisely, for any machine learning model $g_{\theta}\colon D\to\mathbb{R}$ with parameter $\theta$ in a set $\Theta$, $f^{*}$ can be approximated on the set $D\subseteq\mathbb{R}^{d}$ by minimizing the empirical squared error
@@ 3 A posteriori approximation error estimates
Thanks to the implicit form of the Legendre transform (2.1) it is possible to estimate the $L^{2}$-approximation error of any candidate approximator $g\colon D\to\mathbb{R}$ of $f^{*}\colon D\to\mathbb{R}$ with a standard Monte Carlo average (even when $f^{*}$ is not known). Let $\mathcal{X}_{\rm test}$ be a test set of random points $x\in C$ independently sampled from a probability distribution $\mu$ over $C$. Then, due to (2.1), one has
@@
with respect to the push-forward $\nu=\mu\circ(\nabla f)^{-1}$ of the measure $\mu$. This result lets us always report the exact Monte-Carlo estimates of the error $\left\|g-f^{*}\right\|^{2}_{L^{2}(D,\nu)}$, as if $f^{*}$ were known to us in closed form. We further extend this result by adding estimator confidence intervals in Lemma 1 of Appendix A.
@@ 5 Numerical experiments
In this section, we report results of numerical approximations [^6] All computations were performed on NVIDIA T4 GPU and Intel Core Xeon CPUs running Python 3.11.12. More experiments, including the algorithms by Lucet’s and Peyré, are available in the supplementary material. of the convex conjugates of differentiable convex functions using Deep Legendre Transform (DLT) with MLPs, ResNets, ICNNs, and KANs. As an application, we use DLT to approximate solutions of the Hamilton–Jacobi PDE.
5.1 Numerical approximations with MLPs, ResNets, ICNNs
We first approximate $f^{*}\colon D\to\mathbb{R}$ with various neural network architectures, including standard multi-layer perceptrons (MLPs), residual networks (ResNets), and input convex neural networks (ICNNs). MLPs use two hidden layers with 128 units and GELU activation. ResNets apply GELU in two residual blocks, each containing two dense layers, resulting in a 4-layer architecture with skip connections. MLP_ICNNs use the same layer structure as MLPs but constrain hidden weights to be non-negative and apply softplus activations to ensure convexity. ICNNs follow Amos2017, using softplus and non-negative weights in the two hidden layers, along with input skip connections for convexity. We initialized the parameters with Gaussian initialization, but for ICNN, the weights are squared. For training, we used the Adam optimizer777see Adam.
To validate our Deep Legendre Transform (DLT) approach, we benchmark it against direct learning of the convex conjugate (i.e., trained on pairs $(y,f^{*}(y))$) in cases where $f^{*}$ is known analytically. This comparison is crucial, as it establishes an upper bound on achievable performance. We compare our implicit method that trains $g_{\theta}$ to satisfy $g_{\theta}(\nabla f(x))\approx\langle x,\nabla f(x)\rangle-f(x)$ against direct learning that trains the same NN to approximate $f^{*}$ directly as a target function.
Table 1: Function expressions, domains, Legendre transforms, and dual images $D=\nabla f(C)$.

### v2

@@ Abstract
We introduce a novel deep learning algorithm for computing convex conjugates of differentiable convex functions, a fundamental operation in convex analysis with various applications in different fields such as optimization, control theory, physics and economics. While traditional numerical methods suffer from the curse of dimensionality and become computationally intractable in high dimensions, more recent neural network-based approaches scale better, but have mostly been studied with the aim of solving optimal transport problems and require the solution of complicated optimization or max-min problems. Using an implicit Fenchel formulation of convex conjugation, our approach facilitates an efficient gradient-based framework for the minimization of approximation errors and, as a byproduct, also provides a posteriori estimates of the approximation accuracy. Numerical experiments demonstrate our method’s ability to deliver accurate results across different high-dimensional examples. Moreover, by employing symbolic regression with Kolmogorov–Arnold networks, it is able to obtain the exact convex conjugates of specific convex functions. Code is available at https://github.com/lexmar07/Deep-Legendre-Transform
@@ 1 Introduction
In this paper, we introduce DLT, a deep learning method for computing convex conjugates, which play an important role in different fields such as convex analysis, optimization, control theory, physics, and economics.
The modern formulation of convex conjugation, also called Legendre–Fenchel transformation, was introduced by Fenchel (1949)  [^1] The Legendre transform—a fundamental operation for switching between dual formulations and variables—has broad utility across thermodynamics, mechanics, optimization, and economics. Key applications include: switching between thermodynamic potentials by exchanging extensive variables (entropy, volume) for intensive conjugates (temperature, pressure) in physics; constructing Moreau envelopes in variational analysis; deriving indirect utility and profit functions in economics; and computing convex potentials in optimal transport. Recent work has further demonstrated applications in approximate dynamic programming Sharifi Kolarijani and Mohajerin Esfahani (2023) and self-concordant smoothing Adeoye and Bemporad (2023). This paper introduces the Deep Legendre Transform algorithm, establishes its theoretical foundations, and demonstrates its application to optimal-control problems governed by Hamilton–Jacobi equations. and defines the convex conjugate of a function $f\colon C\to\mathbb{R}$ on a nonempty subset $C\subseteq\mathbb{R}^{d}$ by
@@
for all $y\in D=\nabla f(C)$. The right side of (1.2), called Legendre transform, was used by Legendre (1787) when studying the minimal surface problem. Note that for it to be well-defined, the gradient map $\nabla f\colon C\to D$ does not have to be invertible. In fact, it can easily be shown that if $f$ is differentiable and convex, the right side of (1.1) is maximized by any $x\in C$ satisfying $\nabla f(x)=y$. So the preimage $(\nabla f)^{-1}(y)$ in (1.2) can be substituted by any point $x\in(\nabla f)^{-1}(y)$.
@@
based on a finite subset $\mathcal{X}\subseteq C$. Of course, for (1.3) to provide a satisfactory approximation to $f^{*}$, the subsets $\mathcal{X}$ and $\mathcal{Y}$ need to be sufficiently dense. Typically, $\mathcal{X}$ and $\mathcal{Y}$ are chosen as grids $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{d}$ and $\mathcal{Y}=\mathcal{Y}_{1}\times\dots\times\mathcal{Y}_{d}$ of similar size. For instance, for regular grids of the form $\mathcal{X}=\mathcal{Y}=I_{N}^{d}=\{\frac{1}{N},\frac{2}{N},\dots,1\}^{d}$, conditions under which the discrete Legendre–Fenchel transform (1.3) converges to $f^{*}$ have been given by Corrias (1996). In this case, a naive computation of (1.3) for all $y\in\mathcal{Y}$, has complexity $O(N^{2d})$. For the case where NN is a power of 22, Brenier (1989) and Noullez and Vergassola (1994) have derived fast algorithms to compute (1.3), similar in spirit to the fast Fourier transform of Cooley and Tukey (1965), with complexity $O((N\log N)^{d})$. By first passing to convex hulls, Lucet (1997) has been able to design an algorithm with complexity $O(|\mathcal{X}|+|\mathcal{Y}|)$, which for $\mathcal{X}=\mathcal{Y}=I^{d}_{N}$ means $O(N^{d})$. But since this is still exponential in $d$, it essentially limits the method to dimensions $d=1,2,3$ and $4.$
A different approach, going back to a tweet by Peyré (2020), uses the fact that if $C\subseteq\mathbb{R}^{d}$ is bounded, $f^{*}$ can be approximated [^2] The entropic log-sum-exp smoothing of a max can be traced back to the Gibbs variational principle; brought into modern convex-optimisation analysis by Nesterov (2005). with the softmax convex conjugate
@@
Approximating the convolution $Q_{\varepsilon}(f)*G_{\varepsilon}$ with a sum over a finite set $\mathcal{X}\subseteq C$ yields a GPU-friendly algorithm with complexity $O(|\mathcal{X}|\log|\mathcal{X}|)$; see Blondel and Roulet (2024) for details. But since it is grid-based, it still only gives accurate results in low-dimensional settings.
Lucet et al. (2009) have explored the computation of convex conjugates of bivariate piecewise linear quadratic convex functions, and Haque and Lucet (2018) have proposed an algorithm for such functions with linear worst-case complexity. But the scalability to higher dimensions remains a challenge.
Recently, numerical methods for convex conjugation have also been studied in the machine learning literature as a tool to solve the Kantorovich dual of optimal transport problems and develop efficient Wasserstein GANs. While their performance has mostly been tested to see how well they help to solve transport problems, promising results have been obtained by training neural networks to predict an optimal $x$ in the Legendre–Fenchel transform (1.1). E.g. Dam et al. (2019) and Taghvaei and Jalali (2019) predict $x$ from $y$ with a neural network $h_{\vartheta}(y)$, where $h_{\vartheta}$ is maximizing the right side of (1.1). Makkuva et al. (2020) do the same using predictors of the form $h_{\vartheta}=\nabla\ell_{\vartheta}$ for an ICNN [^3] Input Convex Neural Network; see Amos et al. (2017) $\ell_{\vartheta}$. Korotin et al. (2021a, b) and Amos (2023) have proposed to predict an optimal $x$ in (1.1) with a neural network $h_{\vartheta}$ by minimizing the first order condition
@@
$g_{\theta}(\nabla f(x))\approx\left<x,\nabla f(x)\right>-f(x)\quad\mbox{for all }x\in C.$
DLT implements this by solving
@@
    If $g_{\theta}$ is chosen as a neural network, the approach readily scales to high-dimensional setups, and if combined with an ICNN as approximator, it is guaranteed to output a convex approximation to the true Legendre transform, which is known to be convex.
@@
    Specifying $g_{\theta}$ as a KAN [^4] Kolmogorov–Arnold Network; see Liu et al. (2025) makes it possible to obtain exact solutions for certain convex functions.
@@
(i) is a theoretical advantage compared to existing approaches and, as illustrated in Section 5 below, yields good numerical results, in fact, limited only by the used NN architecture/computing power available. We demonstrate that DLT achieves state-of-the-art performance by comparing it to the direct learning of the convex conjugate (i.e., using $f^{*}$ as a target function in cases where it is available in closed form).
(ii) ICNNs have been introduced by Amos et al. (2017), and it has been shown by Chen et al. (2019) that they can approximate any convex function arbitrarily well on compacts. With neural networks as approximators, the minimization problem (1.5) can numerically be solved in high dimensions with a stochastic gradient descent algorithm. The choice of ICNNs ensures that the approximation is convex. This regularizes the learning problem and helps against overfitting. Moreover, it is helpful in various downstream tasks, such as optimization, control, or optimal transport, where convex conjugation can be utilized as a powerful tool to find optimal solutions. Alternatively, we use ResNet (He et al. (2016)), which in many multidimensional cases provides a better approximation than ICNN but does not guarantee convexity.
(iii) KANs have recently been introduced by Liu et al. (2025). They are particularly powerful for approximation tasks requiring high precision. But in addition, they can also be used for symbolic regression to recover functions that have a closed form expressions in terms of special functions such as polynomials, $\exp$, $\log$, $\sin$ or $\cos$. In Section 5.5 below, we show that KANs can compute Legendre transforms of certain convex functions exactly.
(iv) As a byproduct, our approach yields a posteriori $L^{2}$-approximation guarantees. Indeed, by testing a candidate approximating function $g\colon C\to\mathbb{R}$ against the implicit Legendre transform (1.4) on a test set $\mathcal{X}_{\rm test}\subseteq C$ sampled from a distribution $\mu$, it is possible to derive estimates of the $L^{2}$-approximation error $\left\|g-f^{*}\right\|^{2}_{L^{2}(D,\nu)}$ with respect to the push-forward $\nu=\mu\circ(\nabla f)^{-1}$ of $\mu$ under the gradient map $\nabla f\colon C\to D$. This is valuable beyond the framework of this paper since it makes it possible to evaluate the performance of any numerical method for convex conjugation.
@@
The remainder of the paper is organized as follows: In Section 2, we introduce DLT. In Section 3, we derive estimates of the $L^{2}$-approximation error of any candidate approximating function of $f^{*}$ with respect to a given probability measure $\nu$ on the image $D=\nabla f(C)$ of $C$ under the gradient map $\nabla f$. In Section 4, we learn a generalized inverse mapping of $\nabla f$ that improves the performance of DLT if $\nabla f$ is heavily distorting $C$ and is needed to estimate the $L^{2}$-approximation error with respect to a desired target distribution $\nu$ on $D$. In Section 5, we report the results of different numerical experiments. Sec 6 concludes. Proofs of theoretical results and additional numerical experiments are given in the Appendix.
@@ 2 Convex conjugation as a machine learning problem
Let $f\colon C\to\mathbb{R}$ be a differentiable convex function defined on an open convex subset $C\subseteq\mathbb{R}^{d}$. Our goal is to approximate the convex conjugate $f^{*}\colon\mathbb{R}^{d}\to(-\infty,\infty]$ on the image $D=\nabla f(C)$ of $C$ under the gradient map $\nabla f$.
By the Fenchel–Young inequality555see e.g., Fenchel (1949) or Rockafellar (1970), one has
@@
which makes it possible to formulate the numerical approximation of $f^{*}$ as a machine learning problem. More precisely, for any machine learning model $g_{\theta}\colon D\to\mathbb{R}$ with parameter $\theta$ in a set $\Theta$, DLT approximates $f^{*}$ on $D\subseteq\mathbb{R}^{d}$ by minimizing the empirical squared error
@@ 3 A posteriori approximation error estimates
Thanks to the implicit form of the Legendre transform (2.1) it is possible to estimate the $L^{2}$-approximation error of any candidate approximator $g\colon D\to\mathbb{R}$ of $f^{*}\colon D\to\mathbb{R}$ with a standard Monte Carlo average even if the true $f^{*}$ is not known. Let $\mathcal{X}_{\rm test}$ be a test set of random points $x\in C$ independently sampled from a probability distribution $\mu$ over $C$. Then, due to (2.1), one has
@@
with respect to the push-forward $\nu=\mu\circ(\nabla f)^{-1}$ of the measure $\mu$. This result lets us report Monte Carlo estimates [^6] Confidence intervals of the estimator are provided in Lemma 1 in Appendix A. of the error $\left\|g-f^{*}\right\|^{2}_{L^{2}(D,\nu)}$ as if $f^{*}$ were known in closed form.
@@ 5 Numerical experiments
In this section, we report results of numerical approximations [^7] All computations were performed on NVIDIA T4 GPU and Intel Core Xeon CPUs running Python 3.11.12. More experiments, including the algorithms by Lucet’s and Peyré, are available in the supplementary material. of the convex conjugates of differentiable convex functions using DLT with MLPs, ResNets, ICNNs, and KANs. As an application, we use DLT to approximate solutions of Hamilton–Jacobi PDEs.
5.1 Numerical approximations with MLPs, ResNets and ICNNs
We first approximate $f^{*}\colon D\to\mathbb{R}$ with different neural networks such as standard multi-layer perceptrons (MLPs), residual networks (ResNets) and input convex neural networks (ICNNs). We use MLPs with two hidden layers containing 128 units and GELU activation. Our ResNets have two residual blocks, each containing two dense layers, resulting in a 4-layer architecture with skip connections and GELU activation. The employed MLP-ICNNs have the same architecture as the MLPs but constrain hidden weights to be non-negative and apply softplus activations to ensure convexity. The ICNNs are specified as in Amos et al. (2017), using softplus activation and non-negative weights in two hidden layers along with input skip connections. Parameters were initialized according to a Gaussian distribution, but for ICNN, the weights were squared to ensure non-negativity. For training, we used the Adam optimizer888see Kingma and Ba (2017).
As a first test, we compare DLT in the three cases of Table 1, where $f^{*}$ is known analytically, to directly learning $f^{*}$ using pairs $(y_{i},f^{*}(y_{i}))$ as targets.
Table 1: Functions $f$, domains, Legendre transforms, and dual images $D=\nabla f(C)$.